# Minkowski tensors for voxelized data: robust, asymptotically unbiased estimators

### v1

@@ 4.1 Description and analysis
Table 2: Results for the 2-dimensional rectangle $K=[-\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$, the spherical shell $S$ with inner radius 1 and outer radius 2 (see Example 4.10) and $R_{a,b}$ from Example 4.11 with $a=(1,2)^{\top}$, $b=(3,5)^{\top}$ by using the estimator from (4.6) and the result from Corollary 4.1. We intersected $K,S,R_{a,b}$ with a grid of resolution $0.0005$ and choose $R=\varepsilon=0.05$ and $a=0.0005$. We took the average of 25 renditions. All values are rounded to the fourth significant digit.
Tensors Values Estimates Tensors Values Estimates
[…]
($\Phi_{1}^{0,2}$ ($K$))2,2 0.2387 0.2438 ($\Phi_{1}^{0,2}$ ($S$))1,1 0.3754 0.3754
[…]
($\Phi_{1}^{0,2}$ ($R_{a,b}$))2,2 0.3173 0.3178
@@
In the following, we explain the algorithm that we will focus on for the remainder of the paper. The algorithm is based on the idea from [21] (see Section 3.2 there) to estimate the Minkowski tensors of a set $K$ with positive reach by first approximating $K$ by a simpler set $K_{0}$ (for instance, $K_{0}$ could be a finite subset of $K$) and by solving a linear system $Ax=b$. The solution $x$ of the system is the $(d+1)$-dimensional vector containing the estimated Minkowski tensors $\hat{\Phi}_{d}^{r,s}(K_{0}),\dots,\hat{\Phi}_{0}^{r,s}(K_{0})$, where $r,s\in{\mathbb{N}}_{0}$. The vector $b$ contains the estimated Voronoi tensors $\hat{\mathcal{V}}_{R_{0}}^{r,s}(K_{0}),\dots,\hat{\mathcal{V}}_{R_{d}}^{r,s}(K_{0})$, for $d+1$ different values $0<R_{0}<\dots<R_{d}$ of the radius, and the matrix $A$ is a Vandermonde-type matrix. It is shown in [21] that the solution of this system yields the exact values of the Minkowski tensors whenever the exact values of the Voronoi tensors are plugged in and $R_{d}$ is smaller than the reach of $K$. Since inverting a matrix is numerically not stable, we replace the matrix inversion by a least squares problem, which allows us to use more than only $d+1$ equations. However, the basic idea of our algorithm remains to estimate the Minkowski tensors of $K$ via the Voronoi tensors of a finite set $K_{0}$. Here $K_{0}$ can be viewed as a finite approximation of a possibly infinite set $K$. In practice $K_{0}$ is usually considered to be a finite sample of points in $K$, while for generating test data one can intersect $K$ with a grid.
@@
Now fix $a>0$, an integer $n\geq d+1$, and $0<R_{1}<\dots<R_{n}$. One of the key ideas of the algorithm is to replace the matrix inversion from [21] by a linear least squares problem. Extending the linear system from [21, (6)], we define the estimators $\widehat{\Phi}_{j}^{r,s}(K_{0},a,R)$, $i\in\{0,\dots,d\}$, as the solution of the linear least squares problem
$\widehat{\mathcal{V}}_{R_{i}}^{r,s}(K_{0},a)=\sum_{k=0}^{d}r!s!\kappa_{k+s}R_{i}^{k+s}\widehat{\Phi}_{d-k}^{r,s}(K_{0},a,R),\qquad i\in\{1,\ldots,n\},$ (4.15)
@@
where $R:=(R_{1},\dots,R_{n})$. Recall that $\kappa_{n}$ denotes the volume of an $n$-dimensional unit ball. Moreover, a symmetric tensor $T\in\mathbb{T}^{r+s}({\mathbb{R}}^{d})$ of rank $r+s$ is determined by the values $T(e_{i_{1}},\dots,e_{i_{r+s}})$ for $1\leq i_{1}\leq\dots\leq i_{r+s}\leq d$ (see (2.1)). The algorithm computes the tensors $\widehat{\Phi}_{j}^{r,s}(K_{0},a,R)$, $i\in\{0,\dots,d\}$, by solving the least squares problem from (4.15) for each choice of $1\leq i_{1}\leq\dots\leq i_{r+s}\leq d$, using standard methods. The following lemma is well known (see [9, Theorems 1.1.5 and 1.1.6] or [20, Theorem 9]).
@@
As a consequence of Lemma 4.7 the least squares problem from (4.15) is continuous, that is whenever the input vector converges to some vector $b\in{\mathbb{R}}^{n}$, the solution will converge to the solution of the least squares problem with input $b$. In particular, if $K$ is a set with positive reach and $R_{n}$ is less then the reach of $K$, then the tensors $\widehat{\Phi}_{j}^{r,s}(K_{0},a,R)$ from (4.15) converge to ${\Phi}_{j}^{r,s}(K)$ as $\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)$ converges to $\mathcal{V}_{R}^{r,s}(K)$, for which Theorem 4.3 provides sufficient conditions. Note that the approximating tensors $\widehat{\Phi}_{j}^{r,s}(K_{0},a,R)$ depend on the choice of the fixed radii $0<R_{1}<\cdots<R_{n}$, but the Minkowski tensors ${\Phi}_{j}^{r,s}(K)$ will be independent of the radii (as they should be) as long as $K$ has reach greater than $R_{n}$. We would like to point out that although the solution to the least squares problem from Lemma 4.7 involves a matrix inversion, the algorithm for solving the problem does not compute an inverse matrix. Lemma 4.7 merely serves to show that the solution is continuous with respect to the input. In our situation we apply Lemma 4.7 for each choice of $1\leq i_{1}\leq\dots\leq i_{r+s}\leq d$ to
@@
to obtain $\widehat{\Phi}_{d-j}^{r,s}(K_{0},a,R)(e_{i_{1}},\dots,e_{i_{r+s}})$ for $j\in\{0,\dots,d\}$. So there are two main tasks involved in the algorithm: To estimate the Voronoi tensors of $K_{0}$ with respect to the $n$ values $R_{1},\dots,R_{n}$ by the estimator in (4.6) and then to solve the system (4.15). Therefore the only randomness involved in the algorithm is contained in the random grid $\eta$.
We now explain the role of the relevant quantities. As an input the algorithm takes a finite set $K_{0}$, the rank parameters $r,s\in{\mathbb{N}}_{0}$ of the Minkowski tensors, the number $n\geq d+1$ of equations as well as the maximal radius $R_{n}$. The smallest radius $R_{1}$ used is the average nearest neighbor distance in $K_{0}$, that is
$av(K_{0}):=\frac{1}{|K_{0}|}\sum_{x\in K_{0}}\min\{\|x-z\|:z\in K_{0}\setminus\{x\}\},$
@@
instead of $R_{n}$. In this case the algorithm computes the distance of the data from the boundary of the observation window in each direction and and chooses the minimum of these values for $R_{n}$. By the distance of $K_{0}$ from the boundary of $W$ in some direction $s\cdot e_{j}$, where $s\in\{-1,1\}$ and $e_{j}$ denotes the $j$-th standard unit vector, we mean
@@
The output of the algorithm consists of a list containing the $d+1$ estimated Minkowski tensors $\widehat{\Phi}_{d}^{r,s}(K_{0},a,R),\dots,\widehat{\Phi}_{0}^{r,s}(K_{0},a,R)$ (in this order), where a tensor is given as a dictionary whose keys are the indices of the individual values. Note that the tensors are symmetric and therefore the algorithm does not estimate all the values since some of them are redundant. For a tuple of indices $(i_{1},\dots,i_{r+s})\in\{1,\ldots,d\}^{r+s}$ with $i_{1}\leq\dots\leq i_{r+s}$ the algorithm estimates the value of the tensor corresponding to these indices, but not those corresponding to permutations of the tuple since they are the same. Table 3 shows a small summary of the input of the algorithm.
Table 3: Input parameters of the algorithm
@@
$R_{n}$ maximal radius considered (we recommend $R_{n}\geq(d+1)R_{1}$)
@@ 4.2 Implementation
The algorithm is implemented in Python, and the code is available via an open-source package [47]. One can apply the algorithm by using the Python functions Voromink and Vorosurf in the file VorominkEstimation.py. Details can be found in the description of the code, which can also be found in [47]. Closely following the above described estimation of the Minkowski tensors, the user chooses some rank parameters $r,s\in{\mathbb{N}}_{0}$. Then the algorithm will return estimators of the $d+1$ Minkowski tensors $\Phi_{0}^{r,s}$$(K),\dots,$$\Phi_{d}^{r,s}$$(K)$, based on the input data $K_{0}$ and the dimension $d$, for a (in general) unknown set $K$. The algorithm also offers the option to rotate the random grid $\eta$ on which the algorithm is based. This option is useful for test data that lies somehow parallel to the axis spanned by the standard vectors. For real data this rotation should, in general, not be necessary and can be skipped to further reduce the run time (especially in high dimensions).
@@
We found the run times of our code to be convenient for practically relevant examples. For example, a single rendition of the simulations for a two-dimensional rectangle (results can be found in Appendix A), containing about 150,000 points, took an average of 30 seconds. In this case, the algorithm was applied for 5 different rank parameter tuples $(r,s)$ (all possible combinations with $r+s\leq 2$), and for each tuple, 50 runs were performed (resulting in a total of 250 runs and a computation time of 1.5 hours). Computation time scales with the number of points of the grid process $\eta$. Therefore the time increases highly with the dimension $d$ and with a refinement of the resolution of the grid process which is taken as the average nearest neighbor distance $av(K_{0})$ of the input data $K_{0}$. According to Theorem 4.3, it would theoretically be better to choose the grid resolution even smaller than $av(K_{0})$. However, in our simulations, we found that this did not significantly improve the results but considerably increased the runtime.
4.3 Choice of parameters
@@
The choices of the parameters $n,R_{n}$ has a big influence on the results of the algorithm. Therefore we share the experience we made while working with the algorithm. For the parameter $n$ we obtained better results for higher values. In principle one can choose $n$ as big as possible, but the distances $R_{i+1}-R_{i}$ between the values of consecutive radii should not be smaller than the average nearest neighbor distance in the test data. We often used the value $n=50$.
@@
The choice of the parameter $R_{n}$ is much more intricate. The theory requires the value of $R_{n}$ to be smaller than the reach of the original set $K$ that is represented by the data $K_{0}$. In practice this value is usually unknown. If the original set is supposed to be convex, $R_{n}$ could in principle be arbitrarily big. In the general case one should choose this parameter depending on an assessment of the data. It is important that $R_{n}$ is not too small. We recommend it to be at least $d+1$ times the average nearest neighbor distance $av(K_{0})$ in the test data $K_{0}$, but from our experiences this value sometimes needs to be increased further. For more advanced applications and requirements, Bayesian optimization can be a helpful tool for obtaining the best possible parameters.

### v2

@@ 4.2 Voronoi-FD algorithm
Table 2: Results for the two-dimensional rectangle $R_{b}$, $b=(3,5)^{\top}$, with side lengths 3 and 5 (see Example 5.1), the spherical shell $S$ with inner radius 1 and outer radius 2 (see Example 5.2) and $R_{p,q}$ from Example 5.3 with $p=(1,2)^{\top}$, $q=(3,5)^{\top}$ by using the Voronoi-FD algorithm. We intersected $R_{b},S,R_{p,q}$ with a grid of resolution $\Delta=0.0005$ and choose $\varepsilon=0.05$ and $a=0.0005$. We took the average of 25 renditions.
Tensors Exact value Voronoi-FD Tensors Exact value Voronoi-FD
($\Phi_{1}^{0,2}$ ($R_{b}$))1,1 0.3979 0.4029 ($\Phi_{1}^{0,2}$ ($R_{b}$))1,2 0 -$10^{-4}$
($\Phi_{1}^{0,2}$ ($R_{b}$))2,2 0.2387 0.2438 ($\Phi_{1}^{0,2}$ (SS))1,1 0.3754 0.3754
($\Phi_{1}^{0,2}$ (SS))1,2 0 $10^{-4}$ ($\Phi_{1}^{0,2}$ (SS))2,2 0.3755 0.3757
($\Phi_{1}^{0,2}$ ($R_{p,q}$))1,1 0.5560 0.5564 ($\Phi_{1}^{0,2}$ ($R_{p,q}$))1,2 0 -$10^{-4}$
($\Phi_{1}^{0,2}$ ($R_{p,q}$))2,2 0.3173 0.3178
4.3 Voronoi-LSQ algorithm
@@
In the following, we explain another algorithm to estimate Minkowski tensors, which is applicable for arbitrary tensor functionals $\Phi^{r,s}_{j}$. It is based on solving a least squares problem, which is why we call it the Voronoi-LSQ algorithm. Importantly, it offers some practical advantages. We focus on the current improvements in comparison to the original algorithm suggested in [32]. The digitized estimator of the volume tensors and especially the least squares fit with respect to the radial dependence strongly improve the robustness of the algorithm. However, its theoretical foundation is only available for sets with positive reach. Nevertheless, in Section 5.2 (see Table 6), we will see that it provides surprisingly accurate results for a set without positive reach. The Voronoi-LSQ algorithm is based on the idea from [32] (see Section 3.2 there) to estimate the Minkowski tensors of a set $K$ with positive reach by first approximating $K$ by a simpler set $K_{0}$ (for instance, $K_{0}$ could be a finite subset of $K$) and by solving a linear system $Ax=b$. The solution $x$ of the system is the $(d+1)$-dimensional vector containing the estimated Minkowski tensors $\hat{\Phi}_{d}^{r,s}(K_{0}),\dots,\hat{\Phi}_{0}^{r,s}(K_{0})$, where $r,s\in{\mathbb{N}}_{0}$. The vector $b$ contains the estimated Voronoi tensors $\hat{\mathcal{V}}_{R_{0}}^{r,s}(K_{0}),\dots,\hat{\mathcal{V}}_{R_{d}}^{r,s}(K_{0})$, for $d+1$ different values $0<R_{0}<\dots<R_{d}$ of the radius, and the matrix $A$ is a Vandermonde-type matrix. It is shown in [32] that the solution of this system yields the exact values of the Minkowski tensors whenever the exact values of the Voronoi tensors are plugged in and $R_{d}$ is smaller than the reach of $K$. Since inverting a matrix is numerically not stable, we replace the matrix inversion by a least squares problem, which allows us to use more than only $d+1$ equations, which further enhances the robustness of the algorithm. However, the basic idea of the Voronoi-LSQ algorithm remains to estimate the Minkowski tensors of $K$ via the Voronoi tensors of a finite set $K_{0}$ by using the estimator from (4.5). Here $K_{0}$ can be viewed as a finite approximation of a possibly infinite set $K$. In practice $K_{0}$ is usually considered to be a finite sample of points in $K$, while for generating test data one can intersect $K$ with a grid.
@@
Now fix $a>0$, an integer $n\geq d+1$, and $0<R_{1}<\dots<R_{n}$. Extending the linear system from [32, (6)], we define the estimators $\widehat{\Phi}_{j}^{r,s}(K_{0},a,{\sf R})$, $i\in\{0,\dots,d\}$, as the solution of the linear least squares problem
$\widehat{\mathcal{V}}_{R_{i}}^{r,s}(K_{0},a)=\sum_{k=0}^{d}r!s!\kappa_{k+s}R_{i}^{k+s}\widehat{\Phi}_{d-k}^{r,s}(K_{0},a,{\sf R}),\qquad i\in\{1,\ldots,n\},$ (4.21)
@@
where ${\sf R}:=(R_{0},\dots,R_{n})$. Recall that $\kappa_{n}$ denotes the volume of an $n$-dimensional unit ball. Moreover, a symmetric tensor $T\in\mathbb{T}^{r+s}({\mathbb{R}}^{d})$ of rank $r+s$ is determined by the values $T(e_{i_{1}},\dots,e_{i_{r+s}})$ for $1\leq i_{1}\leq\dots\leq i_{r+s}\leq d$ (see (2.1)). The Voronoi-LSQ algorithm computes the tensors $\widehat{\Phi}_{j}^{r,s}(K_{0},a,{\sf R})$, $i\in\{0,\dots,d\}$, by solving the least squares problem from (4.21) for each choice of $1\leq i_{1}\leq\dots\leq i_{r+s}\leq d$, using standard methods. The following lemma is well known (see [12, Theorems 1.1.5 and 1.1.6] or [27, Theorem 9]).
@@
Remark 4.13.
As a consequence of Lemma 4.11 the least squares problem from (4.21) is continuous, that is whenever the input vector converges to some vector $y\in{\mathbb{R}}^{n}$, the solution will converge to the solution of the least squares problem with input $y$. In particular, if $K$ is a set with positive reach and $R_{n}$ is less than the reach of $K$, then the tensors $\widehat{\Phi}_{j}^{r,s}(K_{0},a,{\sf R})$ from (4.21) converge to ${\Phi}_{j}^{r,s}(K)$ as $\widehat{\mathcal{V}}_{R_{i}}^{r,s}(K_{0},a)$ converges to $\mathcal{V}_{R_{i}}^{r,s}(K)$, for which Theorem 4.2 provides sufficient conditions.
The approximating tensors $\widehat{\Phi}_{j}^{r,s}(K_{0},a,{\sf R})$ depend on the choice of the fixed radii $0<R_{1}<\cdots<R_{n}$, but the Minkowski tensors ${\Phi}_{j}^{r,s}(K)$ will be independent of the radii (as they should be) as long as $K$ has reach greater than $R_{n}$. We would like to point out that although the solution to the least squares problem from Lemma 4.11 involves a matrix inversion, the Voronoi-LSQ algorithm for solving the problem does not compute an inverse matrix but computes the unconstrained least-squares solution, for example, using an algorithm as in [14]. Lemma 4.11 merely serves to show that the solution is continuous with respect to the input. In our situation we apply Lemma 4.11 for each choice of $1\leq i_{1}\leq\dots\leq i_{r+s}\leq d$ to
@@
to obtain $\widehat{\Phi}_{d-j}^{r,s}(K_{0},a,{\sf R})(e_{i_{1}},\dots,e_{i_{r+s}})$ for $j\in\{0,\dots,d\}$. So there are two main tasks involved in the Voronoi-LSQ algorithm: To estimate the Voronoi tensors of $K_{0}$ with respect to the $n$ values $R_{1},\dots,R_{n}$ by the estimator in (4.5) and then to solve the system (4.21). Therefore the only randomness involved in the algorithm is contained in the random grid $\eta$.
We now explain the role of the relevant quantities. As an input the Voronoi-LSQ algorithm takes a finite set $K_{0}$, the rank parameters $r,s\in{\mathbb{N}}_{0}$ of the Minkowski tensors, the number $n\geq d+1$ of equations as well as the maximal radius $R_{n}$. The smallest radius $R_{1}$ used is the average nearest neighbor distance in $K_{0}$, that is
$av(K_{0}):=\frac{1}{|K_{0}|}\sum_{x\in K_{0}}\min\{\|x-z\|:z\in K_{0}\setminus\{x\}\},$ (4.23)
@@
instead of $R_{n}$. In this case the algorithm computes the distance of the data from the boundary of the observation window in each direction and chooses the minimum of these values for $R_{n}$. By the distance of $K_{0}$ from the boundary of $W$ in some direction $s\cdot e_{j}$, where $s\in\{-1,1\}$ and $e_{j}$ denotes the $j$-th standard unit vector, we mean
@@
The output of the Voronoi-LSQ algorithm is a list containing the $d+1$ estimated Minkowski tensors $\widehat{\Phi}_{d}^{r,s}(K_{0},a,{\sf R}),\dots,\widehat{\Phi}_{0}^{r,s}(K_{0},a,{\sf R})$ (in this order), where a tensor is given as a dictionary whose keys are the indices of the individual values. Note that the tensors are symmetric and therefore the algorithm does not estimate all the values since some of them are redundant. For a tuple of indices $(i_{1},\dots,i_{r+s})\in\{1,\ldots,d\}^{r+s}$ with $i_{1}\leq\dots\leq i_{r+s}$ the algorithm estimates the value of the tensor corresponding to these indices, but not those corresponding to permutations of the tuple since they are the same. Table 3 shows a small summary of the input of the Voronoi-LSQ algorithm.
Table 3: Input parameters of the Voronoi-LSQ algorithm
@@
$R_{n}$ maximal radius considered (cf. (4.24))
$a$ resolution of the grid $\eta$ (optional; default: $av(K_{0})$)
@@
$\mathrm{jobs}$ number of cores used for parallelization (optional; default: -1, i.e., all available)
@@ 4.4 Implementation
Both the Voronoi-FD and the Voronoi-LSQ algorithms are implemented in Python, and the code is available via an open-source package [60]. One can apply the algorithm by using the Python functions Voronoi_LSQ and Voronoi_FD in the file VorominkEstimation.py. Details can be found in the description of the code, which can also be found in [60]. Closely following the above described estimation of the Minkowski tensors, the user chooses some rank parameters $r,s\in{\mathbb{N}}_{0}$. Then the Voronoi-LSQ algorithm will return estimators of the $d+1$ Minkowski tensors $\Phi_{0}^{r,s}$$(K),\dots,$$\Phi_{d}^{r,s}$$(K)$, based on the input data $K_{0}$ and the dimension $d$, for a (in general) unknown set $K$. Both algorithms also offer the option to rotate the random grid $\eta$ on which the algorithm is based. This option is useful for test data that lies somehow parallel to the axis spanned by the standard vectors. For real data this rotation should, in general, not be necessary and can be skipped to further reduce the run time (especially in high dimensions).
@@
We found the run times of our code for the Voronoi-LSQ algorithm to be convenient for practically relevant examples. For example, a single rendition of the simulations for a two-dimensional rectangle (results can be found in the supplement), containing about 150,000 points, took an average of 30 seconds on a standard personal computer. In this case, the algorithm was applied for 5 different rank parameter tuples $(r,s)$ (all possible combinations with $r+s\leq 2$), and for each tuple, 50 runs were performed (resulting in a total of 250 runs and a computation time of 1.5 hours). The computation time scales with the number of points of the grid process $\eta$. Moreover, the computation time increases strongly with the dimension $d$ and with a refinement of the resolution of the grid process, which is taken as the average nearest neighbor distance $av(K_{0})$ of the input data $K_{0}$. For large data sets, the computation time can be further decreased by increasing the number of cores used in an efficient parallelization. According to Theorem 4.2, it would theoretically be better to choose the grid resolution even smaller than $av(K_{0})$. However, in our simulations, we found that this did not significantly improve the results but considerably increased the runtime.
4.5 Choice of parameters for the Voronoi-LSQ algorithm
@@
The choices of the parameters $n,R_{n}$ have a major influence on the results of the Voronoi-LSQ algorithm. Therefore, we share the experience we made while working with the algorithm. For the parameter $n$ we obtained better results for higher values. We often used the value $n=50$.
@@
The choice of the parameter $R_{n}$ is much more intricate. The theory requires the value of $R_{n}$ to be smaller than the reach of the original set $K$ that is represented by the data $K_{0}$. In practice, this value is usually unknown. If the original set is supposed to be convex, $R_{n}$ could in principle be arbitrarily big. In the general case, one should choose this parameter depending on an assessment of the data. It is important that $R_{n}$ is not too small. We recommend it to be at least $d+1$ times the average nearest neighbor distance $av(K_{0})$ in the test data $K_{0}$, that is,
$R_{n}\,\geq\,(d+1)\,av(K_{0}).$ (4.24)
From our experience, this value sometimes needs to be increased further and (4.24) should be treated as a lower bound. For more advanced applications and requirements, Bayesian optimization can be a helpful tool for obtaining the best possible parameters.
5 Test simulations for the Voronoi-LSQ algorithm
In this section, we demonstrate the performance of the Voronoi-LSQ algorithm for some test cases. Each test data set $A_{0}$ used in this section represents some simple set $A$, like a sphere or a rectangle. The test data set $A_{0}$ is obtained by intersecting $A$ with a cubic grid with resolution (or lattice constant) $\Delta$. Note that in this case $\operatorname{{\sf N}}(A_{0})=av(A_{0})=\Delta$.
As a first step, we compare the Voronoi-LSQ algorithm to the original algorithm [32] that is based on a matrix inversion and that has been implemented in [18, Section 3.2]. For this purpose, we used the same setup and parameters as in Figure 3.4(c) of [18]. Specifically, we estimate the surface area of a two-dimensional ball of radius 1 using the Voronoi-LSQ algorithm. We repeat this estimation using 20 different values of $R_{n}$; these values are equal to those of $R_{\max}$ from Figure 3.4(c) in [18]. In each case, we use $n=50$. As in [18], we choose $\Delta=0.001$. Our estimated values range between 3.14152 and 3.14147 for the different values of $R_{n}$. These values correspond to an error of about only 53–87% of the previous error produced by the matrix inversion algorithm. In that sense, the accuracy can improve by almost up to a factor of two. The improvement compared to the algorithm introduced in Section 3.3 of [18] is even larger (see Figure 3.4(c) in [18]).
In the next two subsections, we test the Voronoi-LSQ algorithm on simple examples from different classes of sets. In Subsection 5.3, we then perform a convergence analysis to examine the convergence rate of the Voronoi-LSQ algorithm for the examples considered in Subsections 5.1 and 5.2.